OpenStudy (chris215):

PLEASE HELP: Find f(x) and g(x) so the function can be expressed as y = f(g(x)). y=7/x^2+10

OpenStudy (lochana):

is it y = (7/x^2) + 10 ?

OpenStudy (lochana):

well if it's so, f(x) =( 7/x) + 10 , g(x) = x^2